# An Exploratory Case Study of Query Plan Representations

### v1

@@ Abstract.
Case Study, Testing, Query Plan, Unified Representation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Empirical studiesccs: Software and its engineering Specification languagesccs: Information systems Database utilities and tools
1. Introduction
@@
Database Management Systems (DBMSs) are fundamental software systems used to store, retrieve, and run queries on data. They are used in almost every computing device (Website, 2022b, c, a); thus, any bug has potentially severe consequences. Recently, automated testing approaches for DBMSs have gained broad adoption. Query Plan Guidance (QPG(Ba and Rigger, 2023) guides test case generation towards diverse query plans aiming to expose more bugs, Cardinality Estimation Restriction Testing (CERT(Ba and Rigger, 2024) identifies performance bugs by examining query plans, and Mozi (Liang et al., 2024) checks the consistency across query plans of the same query to find incorrect query optimizations.
@@
While these approaches have been effective and have found more than 200 previously unknown and unique bugs, implementing them requires a non-trivial effort, as they require DBMS-specific logic to parse and process query plans. A query plan refers to a series of concrete steps to execute a query written in a declarative language, such as Structured Query Language (SQL). DBMSs expose query plans in various formats (e.g., in textual format), in which case we refer to them as serialized query plans. Unlike query languages, for which widely-used, standardized (Guagliardo and Libkin, 2018), and formalized (Guagliardo and Libkin, 2017) languages exist, such as SQL (Website, 2023d), the formats in which serialized query plans are exposed are non-standardized and DBMS-specific. For example, a predicate in the WHERE clause of SQL corresponds to a concrete step to filter data in the query plans of TiDB (Bousquet et al., 2001), but corresponds to a property of another step to scan tables in the query plans of PostgreSQL (Guliato et al., 2009). We refer to the different ways in which serialized query plans are represented as query plan representations. Considering that hundreds of DBMSs exist, [^1] https://dbdb.io/ lists 1001 DBMSs as of August 2024. implementing the above testing methods requires significant effort as they need to account for differences in query plan representations, thus significantly hindering the effectiveness of the above approaches.
@@
We envision that a unified query plan representation would remove the roadblock to implementing the above testing approaches. In this work, we systematically study query plan representations. We present an exploratory case study (Runeson et al., 2012), which is a method to investigate a phenomenon in depth, including both qualitative and quantitative research methods. We collected documents, source code, and third-party applications of the query plans in nine popular DBMSs across five different data models, and summarized the commonalities and differences of query plan representations. Our study shows that query plan representations are based on three conceptual components: operations, properties, and formats. Based on the study, we designed a unified query plan representation, which allows the representation of all conceptual components we studied from DBMS-specific query plans, so that we can easily apply automated testing methods to multiple targets.
@@
Overall, we believe that our work is significant for the software engineering research community, because we focus on how software developers can develop applications based on serialized query plans. We believe that our unified representation reduces the effort to build applications, including testing, on serialized query plans. We include all study results and the prototype in our supplementary materials. We provide a comprehensive website, which provides supportive additional information including illustrative examples, explanations of studied query plans, and example applications. We make the following contributions:
  • A study of query plan representations and results that are available at https://unifiedqueryplan.github.io/, allowing both practitioners and researchers to study the results.
@@
2. Background
Query optimization. DBMSs consider a variety of potential execution strategies for a query written in a declarative language (Mella et al., 2019), and each execution strategy includes specific steps to execute the query. Query optimization is the process of determining an execution strategy, and typically includes these steps: parsing queries into logical query plans that represent the logical steps to execute the query, including operations such as join, converting them into physical query plans, which represent the executable operations to execute the query, such as hash join, and determining a physical query plan to execute.
Query plans. Query plans, typically short for physical query plans, are organized as a Directed Acyclic Graph (DAG) (Neumann, 2005), in which a step outputs data to another step as input followed by directed edges. As query optimizations are specific to the storage engines and query execution engines, query plans are non-standardized. By executing a query with a specific prefix, such as EXPLAIN for most relational DBMSs that support SQL, we can obtain a serialized query plan in a DBMS-specific query plan representation. In this paper, we study various query plan representations. We did not study logical query plans as they usually are not exposed by DBMSs.
3. Query Plan Case Study
We adopted an exploratory case study as the method to investigate query plan representations. The case study, as an empirical method, is used for investigating a contemporary phenomenon in depth and within its real-world context (van Rosmalen et al., 2014) (the “case”). An exploratory case study is a specific case study that generates new questions, propositions, or hypotheses during the study. In this paper, we chose this method, because we wanted to gain an in-depth understanding of how query plans are represented within mature DBMSs. We followed common guidelines of case study research (Runeson et al., 2012) to design and conduct this study.
3.1. Case Study Design
Table 1. The studied nine DBMSs ranging from various data models, development modes, and release dates.
@@ 3.1. Case Study Design
InfluxDB (Anih et al., 2020) 2.7.0 Time-series 2013 28
MongoDB (Ouyang et al., 2022) 6.0.5 Document 2009 5
MySQL (Egea et al., 2010) 8.0.32 Relational 1995 2
Neo4j (Rabuzin et al., 2022) 5.6.0 Graph 2007 21
PostgreSQL (Guliato et al., 2009) 14.7 Relational 1989 4
SQL Server (Website, 1989) 16.0.4015.1 Relational 1989 3
SQLite (Pawlaszczyk, 2022) 3.41.2 Relational 1990 10
SparkSQL (Meng et al., 2016) 3.3.2 Relational 2014 33
TiDB (Bousquet et al., 2001) 6.5.1 Relational 2016 79
@@
Case selection. The case study of this paper is characterized as single-case (Runeson et al., 2012): the query plan representation is the case, while different DBMSs are the units of the analysis. Table 1 shows the DBMSs that we selected for the study. To conduct a representative and comprehensive study, we made a diverse selection of DBMSs. First, we chose DBMSs of various data models: relational, document, graph, and time-series data models, based on the assumption that the DBMSs of different models might have different query plan representations. The relational data model is the most widely-used one (Codd, 1970), and other non-relational models, which are also called NoSQL models, are widely used for maintaining unstructured or semi-structured data (Strauch et al., 2011). Then, we chose both classic and new-generation DBMSs (Pavlo and Aslett, 2016) ranging from release years from the 1980s to the 2010s. The architectures of the chosen DBMSs include standalone DBMSs and an embedded DBMS, SQLite, which runs in the same process as the application. We included both a commercial DBMS, SQL Server, and open-source DBMSs. Apart from conventional DBMSs, we included an analytics engine for large-scale data processing, SparkSQL, which also optimizes queries. To choose widely-used DBMSs for study, we chose them referring to the DBMS ranking website [^2] https://db-engines.com/en/ranking as of August 2024. as shown in the column Rank. We chose the latest release version of each DBMS.
Data collection. We collected data from multiple sources: documents, source code, officially integrated development environments (IDEs), and third-party applications based on query plans. The use of multiple data sources allowed us to perform data source triangulation (Runeson et al., 2012), that is, we could confirm the study results from the above four different types of data sources. We included the detailed lists of the data sources in the supplementary materials for reference.
Data analysis. For each DBMS, we first examined its official documents describing its query plan representation. If the source code was available, we inspected the relevant source code to gain a better understanding of query plan representations. We also ran official test cases in the official IDEs to observe query plan representations in real-world test cases. For the DBMSs that had open-source query plan applications, we further examined how query plan representations are explained and utilized from a third-party perspective. To answer RQ1, we analyzed the query plan representations from the above data sources qualitatively. To answer RQ2, we qualitatively identified conceptual components in the query plan representations, and quantitatively compared these components. To satisfy observation triangulation (Runeson et al., 2012), one author conducted the study, and another author validated the finding against the raw data.
Listing 1: Query plan examples of PostgreSQL and SQLite in text format. Some properties are truncated by … due to page limitation.
@@
6SELECT t1.c0 FROM t0 INNER JOIN t1 ON t0.c0 = t1.c0 WHERE t0.c0 < 100 GROUP BY t1.c0 UNION SELECT c0 FROM t2 WHERE c0 < 10;
7---------------------PostgreSQL-------------------------------------SQLite---------------@HashAggregate@ (cost=62998.82..63009.32 rows=1050...) ‘--@COMPOUND QUERY@ Group Key: t1.c0--LEFT-MOST SUBQUERY
8 ->Append (cost=27150.40..62996.20 rows=1050 width=4) --SCAN t0
9 ->Group (cost=27150.40..62949.08 rows=200 width=4) --SEARCH t1 USING INDEX(c0)...
10 Group Key: t1.c0 --USE TEMP B-TREE FOR GROUP BY
11 ->Gather Set (cost=27150.40..62948.08 rows=400...) --UNION USING TEMP B-TREE
12 Workers Planned: 2 --SEARCH t2 USING INDEX ...
13 ->Group (cost=26150.38..61901.89 rows=200 width=4)
@@
15 ->Set Join (cost=26150.38..56906.48 rows=1998164 width=4)
@@
17 ->Sort (cost=25970.60..26362.39 rows=156719 width=4)
@@
19 ->Parallel Seq Scan on t0 (cost=0.00..10301.95 rows=156719 width=4)
@@ 20 Filter: (c0 < 100)
21 ->Sort (cost=179.78..186.16 rows=2550 width=4)
@@
23 ->Seq Scan on t1 (cost=0.00..35.50 rows=2550 width=4)
24 ->Bitmap Heap Scan on t2 (cost=10.74..31.37 rows=850 width=4)
@@ 25 Recheck Cond: (c0 < 10)
26 ->Bitmap Index Scan on t2_pkey (cost=0.00..10.53 rows=850 width=0)
@@ 27 Index Cond: (c0 < 10)
28Planning Time: 0.124 ms
3.2. Findings Overview
We found that the studied DBMSs share three conceptual components: operations, properties, and formats. As described in section 2, query plans are considered DAGs. In practice, serialized query plans are usually organized in a tree structure, and each node has a key. The keys typically refer to operations, which are concrete steps executed by DBMSs to retrieve, process, or output data in response to a query, such as Full Table Scan, which refers to the step to scan an entire table. Each operation is associated with zero or multiple properties, which involve operation-related information, such as row, which refers to the estimated number of rows returned. Apart from operation-associated properties, plans also have properties associated with them, such as planning time, which refers to the time to generate the query plan. A special case is the key Filter in the TiDB query plans. It represents the condition that the output data from its child node satisfies, so we deem it as a property instead of an operation. DBMSs typically allow serializing query plans in various formats, such as text, table, JSON, and XML.
1 shows two examples of query plan representations of PostgreSQL and SQLite in a textual format. Lines 1–5 show the SQL statements that create and populate the tables. The function generate_series generates data to populate tables t0 and t2. For the query in line 5, the serialized query plan in PostgreSQL is shown in the left part, and that in SQLite is shown in the right part. The bold texts denote operations, and the non-bold texts denote properties. For example, the operation HashAggregate in line 8 has the properties cost, rows, width, and Group Key. The tree structure is denoted through hierarchical indents, in which the operation with longer indents is a child of the operation with short indents. Although both DBMSs are based on the relational data model, and support a textual format, their query plan representations are significantly different.
Table 2. The number of operations and properties in query plan representations.
@@ 3.2. Findings Overview
DBMS Producer Combinator Join Folder Projector Executor Consumer Sum Card. Cost Conf. Status Sum
@@ 3.3. Operations
Classification. We classified operations into seven categories, as shown in the left part of Table 2. On average, every DBMS defines 48 operations in query plans. Neo4j has the most operations, while InfluxDB has no operations. We noticed that Neo4j requires more operations on nodes and relationships of the graph data model, such as the operation to set node properties (Website, 2024e), and SQLite is a lightweight DBMS with a limited number of operations, such as lacking the operations for creating tables. Overall, NoSQL DBMSs have more operations than relational DBMSs. An exception is InfluxDB, which does not define operations. InfluxDB’s operations are disregarded in query plans due to the limited set of operations supported by the single-entity time-series data.
Terminology. We define the operation categories through the following terms. Recall that data models organize data in entities and attributes, so we represent an entity $E=\{a_{1},a_{2},...,a_{n}\}$, in which $a$ is an attribute, and represent a multiset [^3] A collection of elements in which elements may occur more than once. of entities $D=\{E_{1},E_{2},...,E_{n}\}$. $O_{c}(input)=output$ represents that an operation $O$ of the category $c$ receives input from children operations and produces output to the parent operation. Considering that many operations might apply to varying numbers of children, we chose to model the input from all children as a single input. We give a detailed explanation of each operation category as follows.
Producer. The Producer category consists of the operations that retrieve data from storage or return constants instead of from children’s operations. We denote them as $O_{Producer}()=D_{o}$. The operations in the Producer category are data sources of queries, so they are typically leaf nodes of query plans. For example, in 1, the operation SEARCH in the right part of line 11 represents a full table scan, and Bitmap Heap Scan in line 26 represents a data scan from bitmaps in heap memory. Six of nine DBMSs define more than ten operations in the Producer category, because reading data is usually expensive, and thus reads are customized for different scenarios aiming to improve efficiency. For example, indexes (Guttman, 1984; Lehman and Carey, 1986) can be used to efficiently read data.
Combinator. The Combinator category consists of the operations that change the permutation and combination of entities, such as sort and union, with no changes to attributes. We denote them as $O_{Combinator}(\{E_{i}|E_{i}\in D_{i}\})=\{E_{o}|E_{o}\in D_{o},D_{o}\subseteq D_{i}\}$. In 1, the operation Append of PostgreSQL merge entities from different children operations to a single set by the operations in lines 11 and 26, and is associated with query clause UNION in line 6. To execute a similar functionality in SQLite, the operations COMPOUND QUERY and UNION combine data objects by the operations in the right part of lines 8 and 13, and both operations also belong to Combinator category.
Join. The Join category consists of the operations that generate new entities by recombining attributes. We denote them as $O_{Join}(\{E_{i}|E_{i}\in D_{i}\})=\{E_{o}|\forall a_{o}\in E_{o},a_{o}\in D_{o},D_{o}\subseteq D_{i}\}$. In 1, the operation Set Join in line 17 combines two ordered data from the operations in lines 19 and 23 based on the common fields t0.c0 and t1.c0. MongoDB has no Join operations, because it includes only a single document entity for querying and lacks support for combining data from multiple documents. The higher number of Join operations in Neo4j is because we classified the operations on the edges of the graph data model as belonging to the Join category. In the graph data model, edges establish relationships between nodes, and a broader range of operations can be performed on the edges. For example, in Neo4j, executing the simple query MATCH ()-[r]->() WHERE r.title ENDS WITH ’developer’ RETURN r retrieves the relationships whose properties satisfy r.title ENDS WITH ’developer’. The corresponding query plan is shown in Figure 1. Each line in the table represents an operation and associated properties, and the content outside the table is plan-associated properties. The query plan specifies scanning the relationships, which indicates both nodes (i.e., entities) each relationship connects, so the operation UndirectedRelationshipIndexContainsScan belongs to Join.
Figure 1. An example query plan of the Neo4j operations of the Join category.
@@
Folder. The Folder category consists of the operations that derive new entities from a set of entities. Formally, we denote them as $O_{Folder}(\{E_{i}|E_{i}\in D_{i}\})=\{E_{o}|E_{o}=\psi(D_{o}),D_{o}\subseteq D_{i}\}$, in which $\psi()$ denotes a function, such as $MAX()$. In 1, the operations HashAggregate and Group are in the Folder category and represent data aggregation and grouping, respectively. SQLite does not define operations in the Folder category, but shows similar information in properties together with the operations in the Producer category. The operations in the Folder denote DBMSs’ data transformation capability, so most DBMSs support operations in the Folder category.
@@
Projector. The Projector category consists of the operations that remove attributes from all entities. Formally, we denote them as $O_{Projector}(\{E_{i}|E_{i}\in D_{i}\})=\{E_{o}|E_{o}\subseteq E_{i}\}$. No operation in 1 belongs to this category, and 6 of 9 DBMSs have no operations in this category. We observed that these operations correspond to SELECT clauses, and they are not explicitly denoted in query plans.
@@
Executor. The Executor category consists of the operations that make no change to entities and attributes. We denote them as $O_{Executor}({D_{i}})=\{D_{o}|D_{o}=D_{i}\}$. In 1, the operation Gather Set in line 15 merges the data from the operation Group running in other computing nodes for a distributed architecture. DBMSs define these operations to cater to various designs and goals. For example, PostgreSQL defines the operation MEMORIZE to cache the output from node children into memory to speed up processing. We classified these operations into the Executor category. The average number of operations in the Executor category is 13, and SparkSQL has significantly more operations, 43, in the Executor category than others, because it defines multiple operations to interact with other components, such as the Python library pandas.
@@
Consumer. The Consumer category consists of operations that have no output. We denote them as $O_{Consumer}(D_{i})=\{\}$. Apart from queries, which, in the SQL, are SELECT statements, DBMSs also support other statements, such as CREATE and UPDATE in SQL. DBMSs also expose query plans for these statements, and name them as execution plans for wider usage (Website, 2024b). The operations of the Consumer category usually modify stored data or system variables. For example, SparkSQL uses the operation SetCatalogAndNamespace to control a particular system variable, and we assign these operations to the Consumer category.
3.4. Properties

### v3

# Towards a Unified Query Plan Representation

@@ Abstract
Index Terms:
Case Study, Query Plan, Unified Representation
I Introduction
@@
Database Management Systems (DBMSs) are fundamental software systems used to store, retrieve, and run queries on data. They are used in almost every computing device [1, 2, 3]; thus, any bug has potentially severe consequences. Recently, automated testing approaches for DBMSs have gained broad adoption. Query Plan Guidance (QPG[4] guides test case generation towards diverse query plans aiming to expose more bugs, Cardinality Estimation Restriction Testing (CERT[5] identifies performance bugs by examining query plans, and Mozi [6] checks the consistency across query plans of the same query to find incorrect query optimizations.
@@
While these approaches have been effective and have found more than 200 previously unknown and unique bugs, implementing them requires a non-trivial effort, as they require DBMS-specific logic to parse and process query plans. A query plan refers to a series of concrete steps to execute a query written in a declarative language, such as Structured Query Language (SQL). DBMSs expose query plans in various formats (e.g., in textual format), in which case we refer to them as serialized query plans. Unlike query languages, for which widely-used, standardized [7], and formalized [8] languages exist, such as SQL [9], the formats in which serialized query plans are exposed are non-standardized and DBMS-specific. For example, a predicate in the WHERE clause of SQL corresponds to a concrete step to filter data in the query plans of TiDB [10], but corresponds to a property of another step to scan tables in the query plans of PostgreSQL [11]. We refer to the different ways in which serialized query plans are represented as query plan representations. Considering that hundreds of DBMSs exist, [^1] https://dbdb.io/ lists 1001 DBMSs as of August 2024. implementing the above testing methods requires significant effort as they need to account for differences in query plan representations, thus significantly hindering the effectiveness of the above approaches.
@@
We envision that a unified query plan representation would remove the roadblock to implementing the above testing approaches. In this work, we systematically study query plan representations. We present an exploratory case study [12], which is a method to investigate a phenomenon in depth, including both qualitative and quantitative research methods. We collected documents, source code, and third-party applications of the query plans in nine popular DBMSs across five different data models, and summarized the commonalities and differences of query plan representations. Our study shows that query plan representations are based on three conceptual components: operations, properties, and formats. Based on the study, we designed a unified query plan representation, which allows the representation of all conceptual components we studied from DBMS-specific query plans, so that we can easily apply automated testing methods to multiple targets.
@@
We believe that our unified representation reduces the effort to build applications, including testing, on serialized query plans. We include all study results and the prototype in our supplementary materials. [^2] https://zenodo.org/records/15097905 For ease of access, we also provide a comprehensive website, [^3] https://nus-test.github.io/uplan/ which provides supportive additional information including illustrative examples, explanations of studied query plans, and example applications. We make the following contributions:
  • A study of query plan representations and results that allow both practitioners and researchers to study the results.
@@
II Background
Query optimization. DBMSs consider a variety of potential execution strategies for a query written in a declarative language [13], and each execution strategy includes specific steps to execute the query. Query optimization is the process of determining an execution strategy, and typically includes these steps: parsing queries into logical query plans that represent the logical steps to execute the query, including operations such as join, converting them into physical query plans, which represent the executable operations to execute the query, such as hash join, and determining a physical query plan to execute.
Query plans. Query plans, typically short for physical query plans, are organized as a Directed Acyclic Graph (DAG) [14], in which a step outputs data to another step as input followed by directed edges. As query optimizations are specific to the storage engines and query execution engines, query plans are non-standardized. By executing a query with a specific prefix, such as EXPLAIN for most relational DBMSs that support SQL, we can obtain a serialized query plan in a DBMS-specific query plan representation. In this paper, we study various query plan representations. We did not study logical query plans as they usually are not exposed by DBMSs.
III Query Plan Case Study
We adopted an exploratory case study as the method to investigate query plan representations. The case study, as an empirical method, is used for investigating a contemporary phenomenon in depth and within its real-world context [15] (the “case”). An exploratory case study is a specific case study that generates new questions, propositions, or hypotheses during the study. In this paper, we chose this method, because we wanted to gain an in-depth understanding of how query plans are represented within mature DBMSs. We followed common guidelines of case study research [12] to design and conduct this study.
III-A Case Study Design
TABLE I: The studied nine DBMSs ranging from various data models, development modes, and release dates.
@@ III-A Case Study Design
InfluxDB [16] 2.7.0 Time-series 2013 28
MongoDB [17] 6.0.5 Document 2009 5
MySQL [18] 8.0.32 Relational 1995 2
Neo4j [19] 5.6.0 Graph 2007 21
PostgreSQL [11] 14.7 Relational 1989 4
SQL Server [20] 16.0.4015.1 Relational 1989 3
SQLite [21] 3.41.2 Relational 1990 10
SparkSQL [22] 3.3.2 Relational 2014 33
TiDB [10] 6.5.1 Relational 2016 79
@@
Case selection. The case study of this paper is characterized as single-case [12]: the query plan representation is the case, while different DBMSs are the units of the analysis. Table I shows the DBMSs that we selected for the study. To conduct a representative and comprehensive study, we made a diverse selection of DBMSs. First, we chose DBMSs of various data models: relational, document, graph, and time-series data models, based on the assumption that the DBMSs of different models might have different query plan representations. The relational data model is the most widely-used one [23], and other non-relational models, which are also called NoSQL models, are widely used for maintaining unstructured or semi-structured data [24]. Then, we chose both classic and new-generation DBMSs [25] ranging from release years from the 1980s to the 2010s. The architectures of the chosen DBMSs include standalone DBMSs and an embedded DBMS, SQLite, which runs in the same process as the application. We included both a commercial DBMS, SQL Server, and open-source DBMSs. Apart from conventional DBMSs, we included an analytics engine for large-scale data processing, SparkSQL, which also optimizes queries. To choose widely-used DBMSs for study, we chose them referring to the DBMS ranking website [^4] https://db-engines.com/en/ranking as of August 2024. as shown in the column Rank. We chose the latest release version of each DBMS.
Data collection. We collected data from multiple sources: documents, source code, officially integrated development environments (IDEs), and third-party applications based on query plans. The use of multiple data sources allowed us to perform data source triangulation [12], that is, we could confirm the study results from the above four different types of data sources. We included the detailed lists of the data sources in the supplementary materials for reference.
Data analysis. For each DBMS, we first examined its official documents describing its query plan representation. Where the source code was available, we inspected it to gain a better understanding of query plan representations. We also ran official test cases in the official IDEs to observe query plan representations in real-world test cases. For the DBMSs that had open-source query plan applications, we further examined how query plan representations are explained and utilized from a third-party perspective. To answer RQ1, we analyzed the query plan representations from the above data sources qualitatively. To answer RQ2, we qualitatively identified conceptual components in the query plan representations, and quantitatively compared them. To satisfy observation triangulation [12], one author conducted the study, and another author validated the finding against the raw data.
Listing 1: Query plan examples of PostgreSQL and SQLite in text format. Some properties are truncated by … due to space constraints.
@@
6
7----------------------PostgreSQL-----------------------
8EXPLAIN (SUMMARY TRUE) SELECT t1.c0 FROM t0 INNER JOIN t1 ON t0.c0 = t1.c0 WHERE t0.c0 < 100 GROUP BY t1.c0 UNION SELECT c0 FROM t2 WHERE c0 < 10;
9
10HashAggregate (cost=62998.82..63009.32 rows=1050...)
@@
12 ->Append (cost=27150.40..62996.20 rows=1050 width=4)
13 ->Group (cost=27150.40..62949.08 rows=200 width=4)
14 Group Key: t1.c0
15 ->Gather Set (cost=27150.40..62948.08 rows=400...)
16 Workers Planned: 2
17 ->Group (cost=26150.38..61901.89 rows=200...)
18 Group Key: t1.c0
19 ->Set Join (cost=26150.38..56906.48...)
@@ 16 Workers Planned: 2
21 ->Sort (cost=25970.60..26362.39...)
@@
23 ->Parallel Seq Scan on t0 (cost=0.00...)
@@ 24 Filter: (c0 < 100)
25 ->Sort (cost=179.78..186.16 rows=2550...)
@@
27 ->Seq Scan on t1 (cost=0.00..35.50...)
28 ->Bitmap Heap Scan on t2 (cost=10.74..31.37...)
@@ 29 Recheck Cond: (c0 < 10)
30 ->Bitmap Index Scan on t2_pkey (cost=0.00...)
@@ 31 Index Cond: (c0 < 10)
32Planning Time: 0.124 ms
33
34-------------------------SQLite-------------------------
35EXPLAIN QUERY PLAN SELECT t1.c0 FROM t0 INNER JOIN ...;
36
37--COMPOUND QUERY
38 --@LEFT-MOST SUBQUERY@ --@SCAN@ t0 --@SEARCH@ t1 USING AUTOMATIC COVERING INDEX (c0=?) --USE TEMP B-TREE FOR GROUP BY
39 --UNION USING TEMP B-TREE
40 --SEARCH t2 USING COVERING INDEX ...
III-B Findings Overview
We found that the studied DBMSs share three conceptual components: operations, properties, and formats. As described in Section II, query plans are considered DAGs. In practice, serialized query plans are usually organized in a tree structure, and each node has a key. The keys typically refer to operations, which are concrete steps executed by DBMSs to retrieve, process, or output data in response to a query, such as Full Table Scan, which refers to the step to scan an entire table. Each operation is associated with zero or multiple properties, which involve operation-related information, such as row, which refers to the estimated number of rows returned. Apart from operation-associated properties, plans also have properties associated with them, such as planning time, which refers to the time to generate the query plan. A special case is the key Filter in the TiDB query plans. It represents the condition that the output data from its child node satisfies, so we deem it as a property instead of an operation. DBMSs typically allow serializing query plans in various formats, such as text, table, JSON, and XML.
1 shows two examples of query plan representations of PostgreSQL and SQLite in a textual format. Lines 1–5 show the SQL statements that create and populate the tables. The function generate_series generates data to populate tables t0 and t2. For PostgreSQL, executing the statement in line 8 outputs the serialized query plan as shown in lines 10–32. The bold texts denote operations, and the non-bold texts denote properties. For example, the operation HashAggregate in line 10 has the properties cost, rows, width, and Group Key. The tree structure is denoted through hierarchical indents, in which the operation with longer indents is a child of the operation with short indents. Although both DBMSs are based on the relational data model, and support a textual format, their query plan representations are significantly different.
TABLE II: The number of operations and properties in query plan representations. RA: Relational Algebra.
@@ III-B Findings Overview
DBMS Producer Combinator Join Folder Projector Executor Consumer Sum Cardinality Cost Configuration Status Sum
[…]
@@ III-C Operations
[…]
Producer. The Producer category consists of the operations that retrieve data from storage or return constants instead of from children’s operations. These operations realize the algebra operator $\sigma$, which selects data that satisfies a given predicate. The operations in the Producer category are data sources of queries, so they are typically leaf nodes of query plans. For example, in 1, the operation SEARCH in line 43 represents a full table scan, and Bitmap Heap Scan in line 28 represents a data scan from bitmaps in heap memory. Six of nine DBMSs define more than ten operations in the Producer category, because reading data is usually expensive, and thus reads are customized for different scenarios aiming to improve efficiency. For example, indexes [28, 29] can be used to efficiently read data.
Combinator. The Combinator category consists of the operations that change the permutation and combination of tuples, such as sort and union, with no changes to attributes. These operations realize the algebra operators $\cup$, $\cap$, and --. In 1, the operation Append of PostgreSQL merge tuples from different children operations to a single set by the operations in lines 13 and 28, and is associated with query clause UNION in line 8. To execute a similar functionality in SQLite, the operations COMPOUND QUERY and UNION combine data objects by the operations in lines 37 and 42, and both operations also belong to Combinator category.
@@
Figure 1: An example query plan of the Neo4j operations of the Join category.
[…]
@@
Folder. The Folder category consists of the operations that derive new tuples from a set of tuples. These operations realize the algebra operator $\gamma$. In 1, the operations HashAggregate and Group are in the Folder category and represent data aggregation and grouping, respectively. SQLite does not define operations in the Folder category, but shows similar information in properties together with the operations in the Producer category. The operations in the Folder denote DBMSs’ data transformation capability, so most DBMSs support operations in the Folder category.
@@
Projector. The Projector category consists of the operations that remove attributes from all tuples. These operations realize the algebra operator $\Pi$. No operation in 1 belongs to this category, and 6 of 9 DBMSs have no operations in this category. We observed that these operations correspond to column lists in SELECT statements, and they are not explicitly denoted in query plans.
@@
Executor. The Executor category consists of the operations that make no change to tuples and attributes. These operations are typically DBMS-specific internal operations, and do not have a clear correspondence with any algebra operators. In 1, the operation Gather Set in line 15 merges the data from the operation Group running in other processes for parallel execution. DBMSs define these operations to cater to various designs and goals. For example, PostgreSQL defines the operation MEMORIZE to cache the output from node children into memory to speed up processing. We classified these operations into the Executor category. The average number of operations in the Executor category is 13, and SparkSQL has significantly more operations, 43, in the Executor category than others, because it defines multiple operations to interact with other components, such as the Python library pandas.
Distributed DBMSs. Within our studied DBMSs, MongoDB, SparkSQL, and TiDB are distributed DBMSs, which allow executing query plans in parallel across multiple computing nodes [30]. Although they have significantly different execution steps than single-node DBMSs, their query plans have no structural difference from other DBMSs’ query plans. The operations that can be executed in parallel are not explicitly stated in query plans. A major difference is that distributed DBMSs define operations to exchange data across nodes. For example, TiDB defines the operations ExchangeReceiver, ExchangeSender, and Shuffle to send, receive, and shuffle data across nodes. We classified them into Executor. InfluxDB and Neo4j only support the distributed architecture in enterprise editions, while we studied their community editions.
@@
Consumer. The Consumer category consists of operations that have no output. These operations correspond to non-query SQL statements, such as UPDATE, so they lack a counterpart in relational algebra. Apart from queries, which, in SQL, are SELECT statements, DBMSs also support other statements, such as CREATE and UPDATE in SQL. DBMSs also expose query plans for these statements, and name them as execution plans for wider usage [31]. The operations of the Consumer category usually modify stored data or system variables. For example, SparkSQL uses the operation SetCatalogAndNamespace to control a particular system variable, and we assign these operations to the Consumer category.
III-D Properties